Grade 12General Physics

what is jerk...?

raju p
10 Years agoGrade 12

3 Answers

raju p
10 Years ago
the rae of change in accelaration is called jerk
KUNCHAM SAMPATH
10 Years ago
 Jerk is the rate of change of acceleration; that is, the derivative of acceleration with respect to time, and as such the second derivative of velocity, or the third derivative of position. Jerk is defined by any of the following equivalent expressions:\vec j(t)=\frac {\mathrm{d} \vec a(t)} {\mathrm{d}t}=\dot {\vec a}(t)=\frac {\mathrm{d}^2 \vec v(t)} {\mathrm{d}t^2}=\ddot{\vec v}(t)=\frac {\mathrm{d}^3 \vec r(t)} {\mathrm{d}t^3}=\overset{...}{\vec r}(t)
SHANMUKESHWAR
10 Years ago
sudden movement